# Iteration Sums of The Euler’s Totient Function Regarding Powers of Fermat Primes

### v1

@@ Abstract
Euler’s Totient function, a cornerstone of number theory, has been extensively studied and applied across disciplines. Exhibiting amazing patterns, the iterations of the Totient function is the major exploration of this paper. This paper first covers the foundational definitions and well established theorems regarding the Totient function, then builds upon those results to investigate applying the Totient function multiple times, such as $\phi(\phi(\phi(n)))$, which are referred to as iterations. Theorems regarding the end behavior of such iterations is presented. Then, an innovative approach of summation is applied to the iterations of the Totient function, which is in the form of $\phi(n)+\phi(\phi(n))+\phi(\phi(\phi(n)))+\cdots$ that could also be expressed as $\sum\phi^{i}(n)$. Novel theorems regarding the sum are proved for all powers of Fermat Primes, as well as an elegant result for powers of three. This paper provides an initiation to the investigations on the sum of iterated Totient function values.
@@ 1 Introduction
Primes, Divisibility, Factorization, and Modular Arithmetic have long been the central topics of number theory. As one of the oldest branches of mathematics, studies on these concepts have begun from ancient times in Egypt and Greece[15] up to modern times, with numerous scholars investing their time and genius. Many problems were solved, while more problems were proposed. Assisting with proofs, functions have been developed as a major tool in mathematics. The Euler’s Totient function, intertwined with multiple concepts of number theory, is the major concern of this paper. The word “Totient” originates from the Latin word “tot”, which means “so many”, as an answer to the question “Quot”(“How many”)[18]. In other literature, the Toitient function is also referred to as the Phi function, because it is denoted by $\phi(n)$. The Totient function counts the number of positive integers relatively prime to $n$ that are less than or equal to $n$.
@@
When there exists a function, then naturally, the question of applying the function more than once comes to mind. When you iterate a function multiple times, each time the previous output of the function becomes the new input of the function, which again produces an output that could be used for the next iteration. Therefore, the Totient function is a nice function to iterate, given that both its inputs and outputs are positive integers. Every output in the range would be a valid input in the domain of the Totient function, which guarantees that undefined results are not a concern with iterations of the Totient function. This paper provides an innovative approach by summing up iterations of the Totient function, such as $\phi(n)+\phi(\phi(n))+\phi(\phi(\phi(n)))+\cdots$. With this approach, patterns were observed, turning into theorems and conjectures. An original theorem proved was a formula that gives the iterated Totient sums of all powers of Fermat primes, which produced a clean corollary for all powers of 3.
@@
These investigations are important because, first of all, the study of the Totient function impacts various disciplines in mathematics. For example, multiplicative functions are introduced when the Totient function is proved to be multiplicative, but not completely multiplicative. The Chinese remainder theorem plays a role in the proof of The Totient function’s multiplicativity, also affecting many studies of modular arithmetic, such as linear congruences. Then, the iteration sums of the Totient function bring the topic of Fermat Primes, a special type of number with unsolved conjectures. In fact, the Totient function itself has many well know conjecture that still aren’t proven, such as Carmichael’s Conjecture[2] and Lehmer’s Problem[3]. Although this paper focuses on theory instead of application, there are many direct applications of number theory in real life. With regards to the context of this paper, the Chinese Remainder Theorem and the Totient function are utilized in cryptography[4]. Specifically, Euler’s Totient Theorem provided the algorithm for one of the most used public key cryptosystems in the world— RSA[16]. By formalizing patterns into theorems, number theory has laid the foundation for transformative technology advancements.
@@ 2 Historical Background
In 1763, the first definition of the Totient function appeared in Euler’s work, Theoremata arithmetica nova methodo demonstrata[11]. However, Euler did not use a notation or any name for the Totient function, but rather defined it in words. The first notation Euler used for the Totient function appeared in one of his later works in 1784, where he used $\pi N$ to denote $\phi(N)$ and wrote $\pi 1=0$[12]. However, this did not become the commonly used notation nowadays. In 1801, Carl Friedrich Gauss introduced the notation $\phi(n)$, along with the conviction that $\phi(1)=1$, in his number theory book Disquisitiones arithmeticae[1]. Gauss accomplished his book by collecting works of mathematicians such as Euler, Fermat, Euclid, Diophantus, and many others, combining them with his own results in the field of number theory, which created this masterpiece that is still regarded as one of the greatest works of mathematical analysis. Due to the significant impacts of this book, the notation $\phi$ became the well know notation for the Totient function. However, it was not until half a century later that $\phi$ was given the name “Toteint Function”. James Joseph Sylvester, founder of the American Journal of Mathematics, proposed the name “Totient” in his 1879 article On certain ternary cubic-form equations[8]. Three different great mathematicians from different times, collectively established the definition, notation, and name for the Totient function. In addition to its significance in number theory, $\phi(n)$ also showcases the beauty of cumulative contributions in mathematics and the timelessness of mathematical discovery across generations.
The earliest studies of applying the Totient function multiple times, or the iteration of the Totient function, began around 1929 by the works of Pillai[17], which mainly focused on the number of times applying the Totient function to get a number to 1. Studies about this iteration of the Totient function were continued by Erd¨os, Granvilie, Pomerance, and Spiro in 1990[5], as well as many other scholars. Some of the results found by these authors are discussed in the Related Works section.
@@ 3 Definitions and Properties
Here $gcd(a,b)$ denotes the greatest common divisor of $a$ and $b$, which is defined as the largest number in the set of common divisors shared by $a$ and $b$. Note that 1 is relatively prime to 1 by definition because $gcd(1,1)=1$. The concept of relatively prime plays an impotant role in the definiton of the Totient function.
@@
For example, presented below is a table of the first 10 numbers and its Totient function value.
@@
From the 10 values in Table 1, some patterns could be observed and interesting questions arise. Why are most of these $\phi(n)$ even numbers? Why is $\phi(n)$ one less than $n$ for prime numbers? Why does $\phi(2)\phi(5)=\phi(10)$, $\phi(2)\phi(3)=\phi(6)$, but $\phi(2)\phi(4)\neq\phi(8)$ ? The last question will be addressed by multiplicative functions. According to the book An Introduction to Number Theory by Harold Stark[6], the following is the definition for multiplicative functions:
@@
This multiplicative property of functions can be quite useful, because we can break big number into its prime powers and deal with those powers separately. To approach the proof of Euler’s Totient function’s multiplicativity, I will first establish the Linear Congruence Lemma and the Chinese Remainder Theorem.
@@
With the linear congruence established, we can prove the Chinese Remainder Theorem, a Theorem first discovered by the Chinese mathematician Sunzi in the 5th century[19].
@@
In order to divide $m_{1}m_{2}$, $x_{1}$ and $x_{2}$ must have the same remainder when dividing $m_{1}m_{2}$, which is also evident in the below derivation:
@@
The Euler Totient function $\phi(n)$ is a multiplicative function, but not completely multiplicative. First, it is easy to show that it is not completely multiplicative, because $2\cdot 4=8$, but $\phi(2)\phi(4)=2\neq\phi(8)$. Second, it’s multiplicativity can be proved:
@@
For a order pair $(a,b)$ from the set $\mathbb{Z}^{{}^{\prime}}_{m}\times\mathbb{Z}^{{}^{\prime}}_{n}$, there is a corresponding $x$ in $\mathbb{Z}^{{}^{\prime}}_{mn}$ such that
@@
To calculate $\phi(n)$ for large numbers of $n$, it is difficult to count all relatively prime numbers with $n$ like Table 1. Fortunately, it is possible to calculate $\phi(n)$ with the following formula, if its prime factors are known.
@@
As seen above, the left side and the right side of the Corollary statement equals the same expression, completing the proof. ∎
@@ 4 Iterations of the Totient Function
which indicates that the value of the expression decreases at least by 1 every time the Totient function is applied. Then if the Totient function is applied n1n-1 times to $n$, it must decrease to 1, the smallest possible value in the range, because even if it only decrease by 1 each time, n(n1)=1.n-(n-1)=1.
@@
Let $n=5$, the iterations of Totient functions applied to $n$ is demonstrated below,
@@
It is evident that applying the Totient function to the number 5 four times or more would equal 1. Taking this theorem a step further, not only would the iterated Totient function lead any number to 1, but it also must equal 2 at some iteration before it converges to 1. To prove that, the specialty of the numbers 2 and 1 is shown first.
@@
For the simplicity, the problem can be broken into two cases:
@@
    Because $a\in\mathbb{Z}^{+}$, $a>1$ indicates that $a\geq 2$, which secures $2^{a-2}\geq 1$. Now, $m=\phi(n)$ could be used,
@@
    Because $p$ is a prime number other than 2, p1p-1 must be an even number. $k\geq 1$, which secures that $p^{k-1}\geq 1$, along with $\phi(a)\geq 1$. Therefore:
@@
Consider the number $27$. When values of the iterates of the Totient function applied to 27 is recorded:
@@
It is evident that for $m=3$, $\phi(\phi(\phi(27)))=2$. It could be tried for any arbitrary positive integer, and it will always reach the number 2 at some point when the Totient function is iterated enough times. Furthermore, one may notice an interesting property with the above example:
@@
$n=\phi(n)+\phi(\phi(n))+\phi(\phi(\phi(n)))\cdots+\phi(2).$
The iteration is terminated at $\phi(2)$ because it is the first time $\phi(\phi(\cdots\phi(n)\cdots))=1$. Adding more 1s after that would not be quite meaningful.
@@
We proccess by induction on the variable $k$.
@@
    When $k=1$, it could be checked that the statement holds true by simply calculating step by step.
@@
    which could be established by evaluating the left side of the above equation term by term. Theorem 3 tells that
@@
A base case was established, and it is proved that if the statement is true when $k=a$ for any arbitrary number $a\geq 1$, it would imply the correctness of the statement for the next number $k=a+1$, which completes the proof by induction. ∎
One may wonder why is the number 3 so special such that the sums of iterated Totient values of its powers equals its original value. First, 31=23-1=2, so $\phi(3-1)=1$, which “goes away” neatly because 1 is trivial in multiplication. For other prime numbers, if $p-1\neq 2$, then $\phi(p-1)\neq 1$. Therefore, when the phi function is applied again, both p1p-1 and $\phi(p-1)$ would be even by Lemma 2, so the multiplicative property could not be applied because the Totient function is not a completely multiplicative function. However, there is a type of special number for which the general formula on the sums of iterated Totient values is found in this paper – Fermat Primes.
@@
The sequence of Fermat numbers can be found on The On-Line Encyclopedia of Integer Sequences:: A000215[14]. Fermat primes are the prime numbers in this sequence. It was originally conjectured by Fermat that all numbers in this sequence are prime. In 1732, Euler disproved Fermat’s conjecture by showing that $F_{5}=4294967297$ is composite because it is divisible by 641641[9]. The primality of all Fermat Numbers stills remains a puzzle for modern mathematics.
@@
Since all prime numbers in the form of $2^{k}+1$ must be a Fermat prime, $2^{k}+1$ would be used to denote Fermat primes for simplicity.
@@
To evaluate the sum of the above geometric sequence, the “Rolling the Snowball” method is applied. First, add 1 and subtract 1 from the equation, then $1+1=2$, which could be added to 2, so $2+2=4$, and it keeps going on as a snowball rolling bigger and bigger, eventually reaching the last term $2^{k-1}+2^{k-1}=2^{k}$.
@@
which completes the proof regarding the sum of iterated Totient functions on Fermat Primes. ∎
This result complies with the sum of iterated Totient values of $3^{k}$, because $2\cdot 3-3=3$. Similar to powers of three, iterated Totient sums on Fermat Primes could also be generalized to the $k$th power. Using results from Lemma 3, The formula for $p^{2}$ is found to be
@@
The pattern seems obvious, with 2 and 3 being constant coefficients, and the power of $p$ increasing. Therefore, it was soon conjectured that:
@@
We could check to see that the above equation agrees with the proven formula for $3^{k}$, because $2\cdot 3^{k}-3\cdot 3^{k-1}=3^{k}$. Everything points in favor of this formula, but unfortunately, this is not the correct formula. That is one of the reasons why mathematics is beloved. When people observes a pattern, such as Fermat’s Primes conjecture and the above formula, no matter how obvious and appealing the pattern is, it could never be established as the truth before proving it. Modern computers emerge fast, but despite the power to check billions of numbers for a pattern, it is still not regarded as true for all numbers. Rigorous proofs are required to establish formulas, using the same methods ancient Greeks proved their math.
@@
$\phi(p^{k})+\phi(\phi(p^{k}))+\phi(\phi(\phi(p^{k})))\cdots+\phi(2)=\frac{2}{p+1}\left[p^{k}(p-1)+2\left(\frac{p-1}{2}\right)^{k}\right]-1$
@@
    Substituting $k=1$ to the right side of the statment gives
@@
    To achieve this, the left side of the equation will be analyzed term by term. The only prime factor of $p^{a+1}$ is $p$, so by Theorem 3,
@@
    For the following terms, $\phi(\phi(\cdots\phi(p^{a})\cdots))$ will always contain the factor 2, so we can continue applying Corollary 3.1 to make all following terms to the form $\phi(\phi(\cdots\phi(p^{a})\cdots))\cdot\phi(p-1)$, until the term $m+1$, where
@@
    For clarity, the values of the above derivations are recorded in the table below.

### v3

@@ Abstract
Euler’s totient function $\phi(n)$ plays a central role in number theory and is applied in areas such as cryptography. In this paper, we study iterations of the totient function. We first prove that for any integer $n>2$, iteratively applying $\phi$ eventually yields the value $2$. Motivated by this terminal behavior, we examine sums of iterated totient values of the form $\phi(n)+\phi(\phi(n))+\phi(\phi(\phi(n)))+\cdots+\phi(2)$, where the summation terminates at $\phi(2)$. We show that for all integers of the form $n=3^{k}$, this sum is equal to $n$. We then extend this result to all powers of Fermat primes, deriving a closed-form expression for the corresponding summations.
@@ 1 Introduction
Primes, Divisibility, Factorization, and Modular Arithmetic have long been the central topics of number theory. As one of the oldest branches of mathematics, studies on these concepts have begun from ancient times in Egypt and Greece [Number_theory_oldest_branches_in_mathematics] up to modern times, with numerous scholars investing their time and genius. Many problems were solved, while more problems were proposed. Assisting with proofs, functions served as a major tool in mathematics. The Euler’s totient function, intertwined with multiple concepts of number theory, is the major concern of this paper. The word “totient” originates from the Latin word “tot”, which means “so many”, as an answer to the question “Quot”(“How many”) [totient_in_Latin]. In other literature, the Toitient function is also referred to as the Phi function, because it is denoted by $\phi(n)$. The totient function counts the number of positive integers relatively prime to $n$ that are less than or equal to $n$.
@@
When there exists a function, then naturally, the question of applying the function more than once comes to mind. When we iterate a function multiple times, each time the previous output of the function becomes the new input of the function, which again produces an output that could be used for the next iteration. Therefore, the totient function suites well to iterate, given that both its inputs and outputs are positive integers. Every output in the range would be a valid input in the domain of the totient function, ensuring that undefined values never arise. This paper provides a new approach to sum up iterations of the totient function: $\phi(n)+\phi(\phi(n))+\phi(\phi(\phi(n)))+\cdots$. With this approach, we prove a formula that gives the iterated totient sums of all powers of Fermat primes, which produced a clean corollary for all powers of 3.
@@
These investigations are important because, first of all, the study of the totient function impacts various disciplines in mathematics. For example, multiplicative functions are introduced when the totient function is proved to be multiplicative, but not completely multiplicative. The Chinese remainder theorem plays a role in the proof of The totient function’s multiplicativity, also affecting many studies of modular arithmetic, such as linear congruences. Then, the iteration sums of the totient function bring the topic of Fermat Primes, a special type of number with unsolved conjectures. In fact, the totient function itself has many well known conjectures such as Carmichael’s Conjecture [Open_Conjecture_Carmichael] and Lehmer’s Problem [Open_Conjecture_Lehmer]. Although this paper focuses on theory instead of application, there are many direct applications of number theory in real life. With regards to the context of this paper, the Chinese Remainder Theorem and the totient function are utilized in cryptography [chinese_remainder_thm]. Specifically, Euler’s totient Theorem provided the algorithm for one of the most used public key cryptosystems in the world— RSA [RSA]. By formalizing patterns into theorems, number theory has laid the foundation for transformative technology advancements.
@@ 2 Historical Background
In 1763, the first definition of the totient function appeared in Euler’s work, Theoremata arithmetica nova methodo demonstrata [euler_totient]. However, Euler did not use a notation or any name for the totient function, but rather defined it in words. The first notation Euler used for the totient function appeared in one of his later works in 1784, where he used $\pi N$ to denote $\phi(N)$ and wrote $\pi 1=0$[Euler_1775_pi_notation_for_phi], yet this did not become the commonly used notation nowadays. In 1801, Carl Friedrich Gauss introduced the notation $\phi(n)$, along with the conviction that $\phi(1)=1$, in his number theory book Disquisitiones arithmeticae [Gauss_uses_Phi_notation]. Gauss accomplished his book by collecting works of mathematicians such as Euler, Fermat, Euclid, Diophantus, and many others, combining them with his own results in the field of number theory, which created this masterpiece that is still regarded as one of the greatest works of mathematical analysis. Due to the significant impacts of this book, the notation $\phi$ became the well know notation for the totient function. However, it was not until half a century later that $\phi$ was given the name “Toteint Function”. James Joseph Sylvester, founder of the American Journal of Mathematics, proposed the name “totient” in his 1879 article On certain ternary cubic-form equations [name_totient_came_from]. Three different great mathematicians from different times, collectively established the definition, notation, and name for the totient function. In addition to its significance in number theory, $\phi(n)$ also showcases the beauty of cumulative contributions in mathematics and the timelessness of mathematical discovery across generations.
The earliest studies of applying the totient function multiple times, or the iteration of the totient function, began around 1929 by the works of Pillai [Earliest_Studyof_Iteration_on_totient_Function], which mainly focused on the number of times applying the totient function to get a number to 1. Erdös, Granvilie, Pomerance, and Spiro continued studies about this iteration of the totient function in 1990 [termination_of_phi^k]. We discussed some of their results in the Related Works section.
@@ 3 Definitions and Properties
Here $gcd(a,b)$ denotes the greatest common divisor of $a$ and $b$, which is defined as the largest number in the set of common divisors shared by $a$ and $b$. Note that 1 is relatively prime to 1 by definition because $gcd(1,1)=1$.
@@
For example, we present below a table of the first 10 numbers and its totient values.
@@
From the 10 values in Table 1, several interesting questions arise: Why are most of these $\phi(n)$ even numbers? Why is $\phi(n)$ one less than $n$ for prime numbers? Why does $\phi(2)\phi(5)=\phi(10)$, $\phi(2)\phi(3)=\phi(6)$, but $\phi(2)\phi(4)\neq\phi(8)$? The last question will be explained using the concept of multiplicative functions. According to the book An Introduction to Number Theory by Harold Stark [introduction_to_number_theory_book], the following defines multiplicative functions:
@@
This multiplicative property of functions can be quite useful, because we can break big numbers into its prime powers and deal with those powers separately. To approach the proof of Euler’s totient function’s multiplicativity, we will first establish the Linear Congruence Lemma and the Chinese Remainder Theorem.
@@
With the linear congruence established, we can prove the Chinese Remainder Theorem, a Theorem first discovered by the Chinese mathematician Sunzi in the 5th century [sunzi_suanjing].
@@
In order to divide $m_{1}m_{2}$, $x_{1}$ and $x_{2}$ must have the same remainder when dividing $m_{1}m_{2}$, which is also evident in the derivation below:
@@
The Euler totient function $\phi(n)$ is a multiplicative function, but not completely multiplicative. First, it is easy to show that it is not completely multiplicative, because $2\cdot 4=8$, but $\phi(2)\phi(4)=2\neq\phi(8)$. Second, we prove its multiplicativity:
@@
For an order pair $(a,b)$ from the set $\mathbb{Z}^{{}^{\prime}}_{m}\times\mathbb{Z}^{{}^{\prime}}_{n}$, there is a corresponding $x$ in $\mathbb{Z}^{{}^{\prime}}_{mn}$ such that
@@
To calculate $\phi(n)$ for large numbers of $n$, it is difficult to count all relatively prime numbers with $n$ like Table 1. Fortunately, we can calculate $\phi(n)$ with the following formula, if its prime factors are known:
@@
As shown above, the left side and the right side of the Corollary statement equals the same expression, completing the proof. ∎
@@
These two Corollaries offer assistance in proving theorems about iterations of the totient function.
@@ 4 Iterations of the Totient Function
which indicates that the value of the expression decreases at least by 1 every time the totient function is applied. Then if the totient function is applied n1n-1 times to $n$, it must decrease to 1, the smallest possible value in the range, because even if it only decreases by 1 each time, n(n1)=1.n-(n-1)=1.
@@
Let $n=5$, we record the iterations of the totient function applied to $n$ below,
@@
It is evident that applying the totient function to the number 5 four times or more would equal 1. Taking this theorem a step further, not only would the iterated totient function lead any number to 1, but it also must equal 2 at some iteration before it converges to 1. To prove that, we first show the specialty of the numbers 2 and 1.
@@
For simplicity, the problem can be broken into two cases:
@@
    Because $a\in\mathbb{Z}^{+}$, $a>1$ indicates that $a\geq 2$, which ensures $2^{a-2}\geq 1$. Now, $m=\phi(n)$ could be used,
@@
    Because $p$ is a prime number other than 2, p1p-1 must be an even number. $k\geq 1$, which ensures that $p^{k-1}\geq 1$, along with $\phi(a)\geq 1$. Therefore:
@@
Consider the number $27$. When we record values of the iterates of the totient function applied to 27,
@@
It is evident that at the third iteration, $\phi(\phi(\phi(27)))=2$. It could be tried for any arbitrary positive integer, and it will always reach the number 2 at some point when the totient function is iterated enough times. Furthermore, we notice an interesting property with the above example:
@@
$n=\phi(n)+\phi(\phi(n))+\phi(\phi(\phi(n)))+\cdots+\phi(2).$
The iteration is terminated at $\phi(2)$ because it is the first time $\phi(\phi(\cdots\phi(n)\cdots))=1$. Adding more 1s after that would not provide more insight. This result follows directly from Theorem 7.
@@
We proceed by induction on the variable $k$.
@@
    When $k=1$, we check that the statement holds true by simply calculating step by step.
@@
    which we establish by evaluating the left side of the above equation term by term. Theorem 3 tells that
@@
We established a base case, and proved that if the statement is true when $k=a$ for any arbitrary number $a\geq 1$, it would imply the correctness of the statement for the next number $k=a+1$, which completes the proof by induction. ∎
One may wonder why is the number 3 so special such that the sums of iterated totient values of its powers equals its original value. First, 31=23-1=2, so $\phi(3-1)=1$, which “goes away” neatly because 1 is trivial in multiplication. For other prime numbers, if $p-1\neq 2$, then $\phi(p-1)\neq 1$. Therefore, when the phi function is applied again, both p1p-1 and $\phi(p-1)$ would be even by Lemma 2, so the multiplicative property could not be applied because the totient function is not a completely multiplicative function. However, there is a type of special number for which we found the general formula on the sums of iterated totient values – Fermat Primes.
@@
The sequence of Fermat numbers can be found on The On-Line Encyclopedia of Integer Sequences:: A000215 [OEIS_sequence]. Fermat primes are the prime numbers in this sequence. It was originally conjectured by Fermat that all numbers in this sequence are prime. In 1732, Euler disproved Fermat’s conjecture by showing that $F_{5}=4294967297$ is composite because it is divisible by $641$ [Euler_disprove_Fermat_Primes]. The primality of all Fermat Numbers still remains a puzzle for modern mathematics.
@@
Since all prime numbers in the form of $2^{k}+1$ must be a Fermat prime, we would use $2^{k}+1$ to denote Fermat primes for simplicity.
@@
To evaluate the sum of the above geometric sequence, we apply the “Rolling the Snowball” method. First, add 1 and subtract 1 from the equation, then $1+1=2$, which could be added to 2, so $2+2=4$, and it keeps going on as a snowball rolling bigger and bigger, eventually reaching the last term $2^{k-1}+2^{k-1}=2^{k}$.
@@
which completes the proof of the sum of iterated totient functions on Fermat Primes. ∎
This result complies with the sum of iterated totient values of $3^{k}$, because $2\cdot 3-3=3$. Similar to powers of three, iterated totient sums on Fermat Primes could also be generalized to the $k$th power. Using results from Lemma 3, we found the formula for $p^{2}$ is:
@@
The pattern seems obvious, with 2 and 3 being constant coefficients, and the power of $p$ increasing. Therefore, we soon conjectured that:
@@
We checked that the above equation agrees with the proven formula for $3^{k}$, because $2\cdot 3^{k}-3\cdot 3^{k-1}=3^{k}$. Everything points in favor of this formula, but unfortunately, it is not the correct formula. When we observe a pattern, such as Fermat’s Primes conjecture and the above formula, no matter how obvious and appealing the pattern is, we could never establish it as the truth before proving it. Modern computers emerge fast, but despite the power to check billions of numbers for a pattern, it is still not regarded as true for all numbers. Rigorous proofs are required to establish formulas, using the same methods ancient Greeks proved their math.
@@
$\phi(p^{k})+\phi(\phi(p^{k}))+\phi(\phi(\phi(p^{k})))+\cdots+\phi(2)=\frac{2}{p+1}\left[p^{k}(p-1)+2\left(\frac{p-1}{2}\right)^{k}\right]-1$
@@
    Substituting $k=1$ to the right side of the statement gives
@@
    To achieve this, we analyze the left side of the equation term by term. The only prime factor of $p^{a+1}$ is $p$, so by Theorem 3,
@@
    For the following terms, $\phi(\phi(\cdots\phi(p^{a})\cdots))$ will always contain the factor 2, so we can continue applying Corollary 3.1 to all following terms,
    $\phi(\underbrace{\phi[\phi(\cdots\phi(}_{\textit{i}}p^{a})\cdots)])=\underbrace{\phi[\phi(\cdots\phi(}_{\textit{i-1}}p^{a})\cdots)]\cdot\phi(p-1),$
    until the term $m+1$, where
@@
    For clarity, we recorded the values of the above derivations in the table below: